Find the x- and y-intercepts of the equation -6x + 3y = -9 The x-intercept is ( ,0). The y-intercept is (0, ).
I missed class the day we learned about this. Could you please show me how
to find them.

Answer by mananth(16946)

-6x + 3y = -9 The x-intercept is ( ,0). The y-intercept is (0, ).

When the line cuts the y axis x=
plug x = 0 in the equation
-6x+3y=-9
0+3y=-9
y=-3
(0,-3)
When the line cuts the X axis Y= 0
-6x+0=-9
x=-9/-6
x=3/2
(3/2, 0)
CHECK